OpenStudy (anonymous):

Pottersville is 51 miles south of Dubuque. Westview is 60 miles west of Dubuque. How far apart are Pottersville and Westview? Round your answer to the nearest mile.

OpenStudy (anonymous):

If you draw the map of the cities, you will see that it makes a right triangle. 51 and 60 are the two shorter sides, and the length that is missing is the hypotenuse, which is the longest side, and also the side you are trying to find the length of. Using the Pythagorean theorem, and a^2 + b^2 = c^2, you can find the hypotenuse.a = 51. b = 60. and c is the unknown. 51^2 + 60^2 = 6201. The square root of 6201 is 78.746428490440123579119583580972, which is rounded to 79. So 79 is your answer.